# Diffusion-based graph generative methods

### v1

@@ II-B Score-Based Generative Models (SGMs)
where represents the weightning function and .
@@ II-C Stochastic Differential Equations (Score SDEs)
| DDPMs | Molecule generation | EDM [31], MDM [49], MiDi [50], GCDM [51], DiGress [35], GeoDiff [41], DiffBP [32], PMDM [52], DiffSBDD [53], TargetDIff [54], DIFFDOCK [55], DiffAb [56], Anand [57], PROTSEED [58], SMCDiff [59], DiffLinker [60], TSDiff [61], SILVR [62], HierDiff [63], D3FG [64],GraDe-IF [65], RINGER [66], GeoLDM [67], DiffMol [68], GFMDiff [69] |
| Motion generation | MotionDiffuse [33], Ren et al. [70], PriorMDM [71], Modiff [34], MoFusion [72], FLAME [42], MLD [73], Alexanderson [74], EDGE [75], Ahn et al. [76], MoDi [77], BiGraphDiff [78], HumanMAC [79], TCD [80], DiffuPose [81], DiffMotion [82], MDM [83] | |
| Others | DPM-GSP [84], DiffSTG [85], HouseDiffusion [86], EDGE [87], Lee et al. [88], SaGess [89], NAP [90], DIFUSCO [91] | |
| SGMs | Molecule generation | Wu et al. [44], ConfGF [45], ColfNet [92], DGSM [46], ProteinSGM [93], Arts et al. [94] |
| Others | EDP-GNN [95], DruM [96], SLD [97] | |
| Score SDEs | Molecule generation | CDGS [47], DiffMD [98], EigenFold [99], JODO [100], NeuralPLexer [101] |
| Others | GraphGDP [36], GSDM [38], NVDiff [37], HGDM [102], DiffusionNAG [103], Diff-POI [104] |
@@ III Applications
Diffusion models not only achieved outstanding performance for images and language data, but also show great potential in the field of special data structure generation. In this section, we will elaborate on applications of diffusion on molecule generation, motion generation and generation of other forms of graph data.
III-A Molecule generation
@@ III-A1 De novo molecule design
De novo molecule design is suppoesed to generate novel, unique, and valid molecule graphs or geometries from scratch. Initially, early attempts built on variational autoencoders (VAEs) and generative adversarial networks (GANs) focus on 2D molecule graphs, whereas recent research shifts attention towards 3D molecule geometries. However, due to the discrete nature of graphs, how to diffuse and sample discrete features and adjacency matrix still remains a major challenge. DiGress [35] is among the few diffusion-based methods that specifically generate 2D molecule graphs. It diffuses graph data with graph edits and learns the reverse process using a graph transformer network. DiGress shows comparable performance to other one-shot generative methods in terms of diversity, novelty and other evaluation metrics. CDGS [47] introduces a hybrid graph noise prediction approach for global and local context extraction together with a ODE solver that promotes efficient graph sampling.
@@
The first diffusion-based method for de novo 3D molecule design is EDM [31]. It operates on the continuous Cartesian coordinates and discrete atomic number in the denoising process. The design of transition kernel follows EGNN [105], which is an E(n) equivariant graph neural network. EDM shows superior performance compared to previous methods in generating medium size molecules. MDM [49] introduces dual equivariant score neural networks, leveraging two SchNets [106] to capture global and local inter-atomic relationships, respectively. Additionally, latent variables are integrated into the diffusion and reverse process to encourage full exploration. Experimental results show that MDM performs better in large-size molecule generation. The model built by Wu et al. [44] employs SGM instead of DDPM. The author propose a handful of energy function for integrating physical and statistical prior knowledge. MiDi [50] design the reverse kernel named rEGNNs and directly predicts the existence of bonds rather than using predetermined rules. This allows for the simultaneous generation of 2D molecule graphs and their corresponding conformers, resulting in enhanced stability of bonds and validity of molecules. GCDM [51] achieves geometry-completion through GCPNET, a graph neural network with similar geometry learning philosophy as ColfNet [92]. JODO [100] is able to jointly generate 2D molecule graphs and 3D conformations. DiffLinker [60] introduces diffusion models to molecular linker design. The architecture of denoising kernel is similar to EDM [31]. SILVR [62] is conditioned on EDM [31] for fragement merging and linker generation. The model is able to explore new chemical space with ideal conditions cheaply. Research by Arts et al. [94] leverage SGMs and force field to learn coarse-grained molecular dynamics. HierDiff [63] first generates fragment representations instead of deterministic fragment, and then construct atom-level 3D structures. Unlike other studies that perform diffusion on feature and position space, GeoLDM [67] maps molecule geometries into latent space and generate samples by decoding latent features predicted by the diffusion model.
III-A2 Conformation design
@@
Conformation generative methods are supposed to generate valid, diverse, and stable molecule geometries for given 2D molecule graphs. ConfGF [45] is one of the first few diffusion-based methods in this area. It proposes to learn gradient fields of atomic coordinates’ log density and regard them as pseudo-forces that act on atoms. With the gradient fields, samples are generated using Langevin dynamics. DGSM [46] argues that previous methods disregard the long-range interactions between non-bonded atoms. During the denoising process, representations and coordinates in the dynamic graph are estimated using MPNN [107]. GeoDiff [41] lays stress on discussing the induced equivariance of Markov chains evolving with equivariant kernels in diffusion process. An equivariant convolutional layer named GFN is designed to be the transition kernel for molecule graph learning. ColfNet [92] comes up with an orthonormal-based local frame for geometries extraction. Unlike other methods that inject noise to coordinates, Torsion Diffusion [108] predicts conformers while only acting on torsion angles, which drastically reduces the sample space. Therefore, an extrinsic-to-intrinsic score model is introduced to convert 3D coordinates to scores in torsional space. Finally, a Boltzmann generator is specifically designed for the chemical system without costly molecular dynamics. DiffMD [98] introduces ScoreSDEs to molecular dynamics along wih a equivariant geometric transformer network. For conformation generation of macrocycle backbones, RINGER [66] is able to effectvely encode ting geometry and handle the cyclic nature of macrocycles and side chains.
@@ III-A3 De novo ligand design
Instead of generating molecules from scratch, ligand design involves designing molecules specifically bind to target receptors. Structure-based drug design (SBDD) aims at predicting the optimal position and oriantation of a small molecule ligand with high affinity and specificity that bind to specific proteins. In the research of DiffBP [32], the researcher mentions that previous research [109, 110, 111] based on auto-regressive models generate ligand in a sequential manner, which leads to several dilemmas. The sequential modeling of molecules is inconsistent to the global interactions of atoms and may lead to early stopping. DiffBP proposes a target-aware diffusion process to model and generate atoms in molecules all at once. Though it surpasses previous methods at some level, the model still suffers from validity of molecules and the drug-likeness of sub-structures. DiffSBDD [53] also performs one-shot molecule generation based on DDPMs and introduces two strategy, protein-conditioned generation and ligand-inpainting generation. Along with the proposed innovation points is a determined binding dataset derived from Binding MOAD [112]. TargetDiff [54] designs a simple SE(3) GNN kernel to connect generative models and binding affinity ranking to evaluate generated samples. Therefore, the model is capable to improve binding affinity prediction in an unsupervised manner. PMDM [52] leverages the dual equivariant score kernel in MDM [49] for pocket-based ligand design and achieves outstanding performance in multiple metrices. D3FG [64] decompose molecules into different components and is capable to generate molecules with realistic structures, drug properties, and binding affinity.
@@ III-A4 Ligand docking
Apart from generating novel ligand, molecular ligand docking plays a essential role in drug discovery by predicting positions and conformations of given molecules bound to target proteins, is also a crucial task in drug discovery. Unlike previous methods based on regression frameworks, DIFFDOCK [55] designs a diffusion model over degrees of freedom, including translational, rotational, and torsional. The implementation of diffusion-based paradigm significantly lifts the success rate of predictions and reduces inference time. EDM-Dock [113] utilizes two EGNNs for extraction of geometric and chemical information of proteins and ligands respectively and a multilayer perceptron for protein-ligand pairwise distance predictions. DPL [114] introduce a novel reverse kernel for their diffusion framework, which consists of input featurization, residual feature update, and equivariant denoising. NeuralPLexer [101] is a framework capable of generating protein-ligand complex. The paper proposes contact prediction module to generate inter-molecular distance distributions, contact maps, and pair representations.
@@ III-A5 Protein design
In the field of protein design, there are three main tasks, which are antigen-specific antibody design, protein sequence and structure design, and motif-scaffold design. Antibody design aims to generate antibodies bind to target antigen. Protein sequences and structure design involves creating and modifying protein sequences and structures to improve stabilities or other properties. Motif-scaffold design scaffoles for desired motifs, a kind of sub-structures that play important roles in the function of proteins. Focusing on the first task, DiffAb [56] generates sequences and structures of complementarity-determining regions (CDR) of antibodies conditioned on antigen structures. Besides, the model is capable to design side-chain orientations of amino acids. Research by anand [57] introduces diffusion-based method to generate protein structures, sequences and rotamers successively. However, this approach may cause inconsistencies and expensive inferences. To tackle this issue, PROTSEED [58] proposes to co-design structures and sequences. The model first extract constraints from a trigonometry-aware context encoder. An equivariant decoder is also designed to update positions of , rotation matrix, and amino acid types. Based on Score-SDE, ProteinSGM [93] generates 6D coordinates of proteins. SMCDiff [59] is proposed to design stable scaffold to support a desired motif. Based on EGNN [105], ProtDiff generates protein backbone in the first place, and SMCDiff later generates scaffolds. For inverse protein folding, GraDe-IF [65] predicts joint distribution of animo acids conditioned on node properties. EigenFold [99] recognize proteins as dynamic structural with conformational flexibility. The model introduces harmonic diffusion and view the reverse process as a cascading-resolution process.
III-B Motion generation

### v2

@@ II-B Score-Based Generative Models (SGMs)
where represents the weighting function and .
@@ II-C Stochastic Differential Equations (Score SDEs)
II-D Brief Summary of Three Paradigms of Diffusion Models
DDPMs, SGMs, and Score SDEs differs from each other in many aspacts. DDPMs model the forward and reverse process using discrete Markov-chain, which provides strong theoretical guarantees. The forward process of DDPMs includes noise injection and signal attenuation, which are both controlled by the schedule . The objective function, KL divergence between ’gronud truth’ noisy samples and parameterized ones, provides better guarantee for convergence. In SGMs, Langevin dynamics are used to estimate the score of samples. Additionally, SGMs can be relatively sensitive to the choice of noise levels compare to the other two classes of diffusion methods. Score SDEs provide a continuous-time stochastic differential equation and regards the generation of sample as the solution of it. Such paradigm is theoretically robust and grounded.
For graph generative models that directly operate on continuous spaces, such as geometric spaces [31, 49, 33] and latent variable spaces [50, 51], they can readily employ the three aforementioned paradigms as their diffusion backbones. However, for those graph generative moethods on discrete spaces, such as graph topologies [35, 52] and discrete feature variables [53], they primarily adopt DDPM as the backbone. This is because the forward and reverse process of DDPM can effectively align with their respective model designs.
II-E Enhancement of Adaptability of Diffusion on Graphs
Due to the anisotropic structure of graphs, adapting diffusion to better align the generation of graphs attracts wide attention.
The most straightforward manner to mitigate the challenge is to introduce diffusion to discrete graph structure. Digress [35] incorporate transition probability matrices into the discrete diffusion process. Thie conditional probability of noisy graph takes the form of:
where . By considering distributions of nodes and edges as Bernoulli distributions, EDGE [52] introduces the discrete diffusion that gradually removes or adds edges during the forward or reverse process. Combined with the efficient MPNN [54], EDGE performs well in generating large graphs while preserving statistical properties of original graphs.
By investigating into the distribution of graph data during the forward process, DDM[55] finds out that adding standard Gaussian noise results in the rapid degradation of signal-to-noise ratios (SNRs), which limits the ability to learn anisotropic structure. Therefore, DDM introduces directional noise, which transform isotropic Gaussian noise into anisotropic noise, to tackle the challenge. Take node faature of noisy graph for an example, it takes the form of:
The customized directional noise ensures the moderate variation of SNRs. To leverage the generation capability of diffusion in concert with the ability of hyperbolic embeddings to capture hierarchical structures, HGDM [56] introduces hyperbolic graph variational auto-encoder for hyperbolic node features learning and two score model for generation in both Euclidean and hyperbolic space. HypDiff [51] addresses anisotropy of non-Euclidean structure for graph latent diffusion. Two geometric constraints are introduced to preserve topologic information as well.
In addition to the aforementioned modifications to diffusion, many studies have chosen to fully leverage the excellent modeling capabilities of diffusion on continuous variables. By constructing proper graph encoders and decoders, a number of studies, including NVDiff [37], Graphusion [57], LDM-3DG [58], HypDiff [51], GeoLDM [50], and LatentDiff [59], perform latent diffusion on implicit representations. In this manner, models exhibit remarkable performances without necessitating complex modifications to diffusion models.
II-F Comparisons and Connections with other generative methods
Before the rise of diffusion, graph generation was primarily based on VAEs, ARs, GANs, and NFs.
Compares to VAEs, diffusion models show better capability of generating high quality samples. However, the denoising process makes it slower for diffusion to generate samples. Compares to ARs, diffusion models can easily generate large samples with high quality. The sequential modeling of ARs leads to relatively weaker ability to model graph with large scale. By modeling samples with different level of noises, diffusion is more stable to train compares to GANs. NFs provide tractable likelihoods and generate samples in an efficient manner. However, the samples generated by NFs are inferior to those generated by diffusion in terms of both validity and fidelity. The performance difference between E-NF and EDM is the best example of this.
The vast majority of models mentioned in this paper are built on a single diffusion model. However, a small subset of them is constructed on the basis of multiple generative models.
With VAEs: Many methods, such as Graphusion, GeoLDM, LatentDiff, and LDM-3DG [57, 50, 59, 58], performs diffusion and denoising process on the latent space and leverage autoencoders to generate samples givin the prior of latent variables. Such paradigm can be seen as an evolution of tradition VAE-based graph generative methods. By combining the effciency and representational power of latent features with the outstanding generative capabilities of diffusion models, these type of methods exhibit great preformance on graph generation tasks.
With ARs: Similar to the aforementioned methods combined with VAE, graph diffusion models combined with autoregressive models are also constructed by building the diffusion model on the foundation of the autoregressive model. For example, AutoDiff [60] and Pard [61] follow the paradigm that first generate global structures with autoregressive model and then refine local structure based on diffusion models. This paradigm helps ensure the reasonableness of local structures in diffusion on large graphs and reduces the computational overhead of diffusion.
@@ II-F Comparisons and Connections with other generative methods
DDPMs | AI for Scientific Discovery | EDM [31], MDM [62], MiDi [53], GCDM [63], DiGress [35], GeoDiff [41], DiffBP [32], PMDM [64], DiffSBDD [65], TargetDIff [66], DIFFDOCK [67], DiffAb [68], Anand [69], TSDiff [70], SILVR [71], SMCDiff [72], DiffLinker [73], PROTSEED [74], HierDiff [75], D3FG [76],GraDe-IF [77], RINGER [78], GeoLDM [50], DiffMol [79], GFMDiff [80], MUDiff [81], DecompDiff [82], AbDiffuser [83], SPDiff [84], LatentDiff [59], DiffCSP [85], DiffCSP++ [86], GemsDiff [87], CrysDiff [88], GHP-MOFassemble [89] |
Computer Vision | MotionDiffuse [33], Ren et al. [90], PriorMDM [91], Modiff [34], MoFusion [92], FLAME [42], MLD [93], Alexanderson [94], EDGE [95], Ahn et al. [96], MoDi [97], BiGraphDiff [98], MDM [99], TCD [100],DiffuPose [101], DiffMotion [102], HumanMAC [103], CommonScenes [104], DiffuScene [105], InstructScene [106], Wei et al. [107] | |
Generic Graph Generation & Others | DPM-GSP [108], DiffSTG [109], HouseDiffusion [110], EDGE [52], EDGE++ [111], Lee et al. [112], NAP [113], SaGess [114], DIFUSCO [115], DDM [55], HypDiff [51], Bergmeister et al. [116], Pard [61] | |
SGMs | AI for Scientific Discovery | Wu et al. [44], ConfGF [45], ColfNet [117], DGSM [46], ProteinSGM [118], Arts et al. [119], VoxMol [120] |
Generic Graph Generation & Others | EDP-GNN [121], DruM [122], SLD [123] | |
Score SDEs | AI for Scientific Discovery | CDGS [47], DiffMD [124], EigenFold [125], JODO [126], NeuralPLexer [127], EEGSDE [128], SD [129], MuDM [130], DiffBindFR [131] |
Computer Vision | DiffuseSG [132] | |
Generic Graph Generation & Others | GraphGDP [36], GSDM [38], NVDiff [37], HGDM [56], DiffusionNAG [133], Diff-POI [134], Graphusion [57] |
@@ III Applications
Diffusion models not only achieved outstanding performance for images and language data, but also show great potential in the field of special data structure generation. In this section, we will elaborate on applications of diffusion on AI for scientific discovery, computer vision, generic graph generation and other forms of graph generative tasks.
III-A AI for Scientific Discovery
@@ III-A1 De novo molecule design
De novo molecule design is suppoesed to generate novel, unique, and valid molecule graphs or geometries from scratch. Initially, early attempts built on variational autoencoders (VAEs) and generative adversarial networks (GANs) focus on 2D molecule graphs, whereas recent research shifts attention towards 3D molecule geometries. However, due to the discrete nature of graphs, how to diffuse and sample discrete features and adjacency matrix still remains a major challenge.
2D Molecule Desgin. DiGress [35] is among the few diffusion-based methods that specifically generate 2D molecule graphs. It diffuses graph data with graph edits and learns the reverse process using a graph transformer network. DiGress shows comparable performance to other one-shot generative methods in terms of diversity, novelty and other evaluation metrics. CDGS [47] introduces a hybrid graph noise prediction approach for global and local context extraction together with a ODE solver that promotes efficient graph sampling.
@@
3D Molecule Design. The first diffusion-based method for de novo 3D molecule design is EDM [31]. It operates on the continuous Cartesian coordinates and discrete atomic number in the denoising process. The design of transition kernel follows EGNN [135], which is an E(n) equivariant graph neural network. EDM shows superior performance compared to previous methods in generating medium size molecules. MDM [62] introduces dual equivariant score neural networks, leveraging two SchNets [136] to capture global and local inter-atomic relationships, respectively. Additionally, latent variables are integrated into the diffusion and reverse process to encourage full exploration. Experimental results show that MDM performs better in large-size molecule generation. The model built by Wu et al. [44] employs SGM instead of DDPM. The author propose a handful of energy function for integrating physical and statistical prior knowledge. GCDM [63] achieves geometry-completion through GCPNET, a graph neural network with similar geometry learning philosophy as ColfNet [117]. JODO [126] is able to jointly generate 2D molecule graphs and 3D conformations. DiffLinker [73] introduces diffusion models to molecular linker design. The architecture of denoising kernel is similar to EDM [31]. SILVR [71] is conditioned on EDM [31] for fragement merging and linker generation. The model is able to explore new chemical space with ideal conditions cheaply. Research by Arts et al. [119] leverage SGMs and force field to learn coarse-grained molecular dynamics. HierDiff [75] first generates fragment representations instead of deterministic fragment, and then construct atom-level 3D structures. Unlike other studies that perform diffusion on feature and position space, GeoLDM [50] and LDM-3DG [58] map molecule geometries into latent space and performs diffusion process in latent space. GFMDiff [80] addresses the challenge of bond generation in 3D molecule generation and builds GFLoss based on pre-defined rules. Such loss term guides the molecule in generating authentic bonds at each sampling step. Instead of point clouds, VoxMol [120] voxelizes molecules in voxel grid, where the length of each cubic grid is .25Å. Such approach does not requires to know the number of atoms that need to be generated and scales better compares to previous methods.
Joint 2D & 3D Molecule Design. Most methods for 3D molecule design operates on 3D point clouds and generate bonds after the sampling phase. Technically speaking, such paradigm is much easier since it does not explicitly generate bonds during the diffusion process. MiDi [53] designs the reverse kernel named rEGNNs and directly predicts the existence of bonds rather than using predetermined rules. This allows for the simultaneous generation of 2D molecule graphs and their corresponding conformers, resulting in enhanced stability of bonds and validity of molecules. MUDiff [81] introduces the transformer-based MUformer, which encodes structural, spectural, and velocity features. Guassian noises and discrete noises are separately added to atom features and edge features.
Conditional Molecule Design. Though most methods mentioned above are designed with the native ability to generate molecule with desired properties, some studies are dedicated to enhancing such capabilities. For those methods that are not specifically designed for this task, the common approach to perform property conditioned molecule generation takes two steps. The first step includes training an graph neural network that predicts certain molecule properties, and the second step is to concatenate desired properties with node features at both training and sampling stages. EEGSDE [128] incorporates an energy guidance function into the stochastic differential equation to guide denoising process. MuDM [130] is a training-free methods, which leverages time-independent posterior approximation and MC sampling. It also achieves remarkable performance in multi-condition generation task by introducing multi-condition guidance derived from probabilistic graphs.
III-A2 Conformation generation and optimization
@@
The precise yet costly molecular dynamics (MD) is crucial for atomistic system simulation. Recent studies predicts 3D molecule conformations with the assistence of deep generative methods with less computational cost. ConfGF [45] is one of the first few diffusion-based methods in this area. It proposes to learn gradient fields of atomic coordinates’ log density and regard them as pseudo-forces that act on atoms. With the gradient fields, samples are generated using Langevin dynamics. DGSM [46] argues that previous methods disregard the long-range interactions between non-bonded atoms. During the denoising process, representations and coordinates in the dynamic graph are estimated using MPNN [54]. GeoDiff [41] lays stress on discussing the induced equivariance of Markov chains evolving with equivariant kernels in diffusion process. An equivariant convolutional layer named GFN is designed to be the transition kernel for molecule graph learning. ColfNet [117] comes up with an orthonormal-based local frame for geometries extraction. Unlike other methods that inject noise to coordinates, Torsion Diffusion [49] predicts conformers while only acting on torsion angles, which drastically reduces the sample space. Therefore, an extrinsic-to-intrinsic score model is introduced to convert 3D coordinates to scores in torsional space. Finally, a Boltzmann generator is specifically designed for the chemical system without costly molecular dynamics. DiffMD [124] introduces ScoreSDEs to molecular dynamics along wih a equivariant geometric transformer network. For conformation generation of macrocycle backbones, RINGER [78] is able to effectvely encode ting geometry and handle the cyclic nature of macrocycles and side chains. DiSCO [137] employs SE(3) Schrödinger bridge to predict atom cordinates and aligns distribution of samples with ground-truth energy landscape.
@@ III-A3 De novo ligand design
Instead of generating molecules from scratch, ligand design involves designing molecules specifically bind to target receptors. Structure-based drug design (SBDD) aims at predicting the optimal position and oriantation of a small molecule ligand with high affinity and specificity that bind to specific proteins. In the research of DiffBP [32], the researcher mentions that previous research [138, 139, 140] based on auto-regressive models generate ligand in a sequential manner, which leads to several dilemmas. The sequential modeling of molecules is inconsistent to the global interactions of atoms and may lead to early stopping. DiffBP proposes a target-aware diffusion process to model and generate atoms in molecules all at once. Though it surpasses previous methods at some level, the model still suffers from validity of molecules and the drug-likeness of sub-structures. DiffSBDD [65] also performs one-shot molecule generation based on DDPMs and introduces two strategy, protein-conditioned generation and ligand-inpainting generation. Along with the proposed innovation points is a determined binding dataset derived from Binding MOAD [141]. TargetDiff [66] designs a simple SE(3) GNN kernel to connect generative models and binding affinity ranking to evaluate generated samples. Therefore, the model is capable to improve binding affinity prediction in an unsupervised manner. PMDM [64] leverages the dual equivariant score kernel in MDM [62] for pocket-based ligand design and achieves outstanding performance in multiple metrices. D3FG [76] decompose molecules into different components and is capable to generate molecules with realistic structures, drug properties, and binding affinity. AutoDiff [60] introduces a diffusion-based autoregressive model, which generate molecules motif-by-motif. Diffusion model is adopted to predict torsional angles between fragments. DecompDiff [82] also address SBDD task by treating atoms in ligand molecules differently. By decomposing molecules into scaffolds and arms and introducing validity guidance, it generates ligand molecules with better drug-likeness and synthesizability.
@@ III-A4 Ligand docking
Apart from generating novel ligand, molecular ligand docking plays a essential role in drug discovery by predicting positions and conformations of given molecules bound to target proteins, is also a crucial task in drug discovery. Unlike previous methods based on regression frameworks, DIFFDOCK [67] designs a diffusion model over degrees of freedom, including translational, rotational, and torsional. The implementation of diffusion-based paradigm significantly lifts the success rate of predictions and reduces inference time. EDM-Dock [142] utilizes two EGNNs for extraction of geometric and chemical information of proteins and ligands respectively and a multilayer perceptron for protein-ligand pairwise distance predictions. DPL [143] introduce a novel reverse kernel for their diffusion framework, which consists of input featurization, residual feature update, and equivariant denoising. NeuralPLexer [127] is a framework capable of generating protein-ligand complex. The paper proposes contact prediction module to generate inter-molecular distance distributions, contact maps, and pair representations. DiffBindFR [131] emphasizes the importance of protein side chian conformation modeling and proposes full-atom graph construction. Pocket side chan torsional score, ligand rotatable bond torsional score, ligand rotational score, and ligand transitional score are introduced to solve the reverse SDE.
@@ III-A5 Protein design
In the field of protein design, there are three main tasks, which are antigen-specific antibody design, protein sequence and structure design, and motif-scaffold design.
Antibody design: Antibody design aims to generate antibodies bind to target antigen. Focusing on this particular task, DiffAb [68] generates sequences and structures of complementarity-determining regions (CDR) of antibodies conditioned on antigen structures. Besides, the model is capable to design side-chain orientations of amino acids. AbDiffuser [83] introduce a novel SE(3) network, which implicitly models residue-to-residue relations and handles sequences with various length better.
@@
Protein sequence and structure design: Protein sequences and structure design involves creating and modifying protein sequences and structures to improve stabilities or other properties. Research by anand [69] introduces diffusion-based method to generate protein structures, sequences and rotamers successively. However, this approach may cause inconsistencies and expensive inferences. To tackle this issue, PROTSEED [74] proposes to co-design structures and sequences. The model first extract constraints from a trigonometry-aware context encoder. An equivariant decoder is also designed to update positions of , rotation matrix, and amino acid types. Based on Score-SDE, ProteinSGM [118] generates 6D coordinates of proteins. LatentDiff [59] first maps 3D protein structures to latent spaces using autoencoders and integrates latent diffusion to protein structure genration.
Motif-scaffold design: Motif-scaffold design scaffoles for desired motifs, a kind of sub-structures that play important roles in the function of proteins. SMCDiff [72] is proposed to design stable scaffold to support a desired motif. Based on EGNN [135], ProtDiff generates protein backbone in the first place, and SMCDiff later generates scaffolds. For inverse protein folding, GraDe-IF [77] predicts joint distribution of animo acids conditioned on node properties. EigenFold [125] recognize proteins as dynamic structural with conformational flexibility. The model introduces harmonic diffusion and view the reverse process as a cascading-resolution process.
III-A6 Other applications
Apart from applications on drug discovery, many studies proposes diffusion-based graph generative methods on many other applications on scientific discovery.
Crystal structure prediction: The goal of crystal structure prediction is to generate stable 3D structures based on compositions of compounds. Traditional methods mainly integrates Density Functional Theory (DFT) [144] to compute the ennergy of compound structures and guide the search to local minima of energy surface. Applications of deep generative methods on this taks could reduce the computational cost. DiffCSP [85] is one of the first methods that focus on this task. It jointly predicts atom fractional coordinates and lattice vectors. On this basis, DiffCSP++ [86] emoloys space group constraint, which is presented in the form of basis constraint of O(3)-invariant logarithmic space of lattice matrix and the Wyckoff position costraint of atom coordinates. GemsDiff [87] introduce a new metric named Frechet ALIGNN Distance (FAD) to provide robust evaluation of generated crystals. CrysDiff [88] design a reconstruction pre-training task and enhance performance on crystal property prediction, which is constrained by limited number of labeled data in previous studies. GHP-MOFassemble [89] presents a method for metal-organic frameworks (MOFs) discovery, which shows great potential in capture. Lee et al. [112] utilizes implicit diffusion moedels for microstructure characterization and reconstruction.
Others: Brain Diffuser [145] shapes the structural brain networks from tensor images and shows better ability in exploiting structutal connectivity features and disease-related information. By adapting the problem of crowd simulation to spatial graph generation, SPDiff [84] designs multi-frame rollout training algorithm to simulate pedestrian trajectories.
III-B Computer Vision
When it comes to the applications of generative graph diffusion models in the field of computer vision (CV), motion generation and scene graph generation are the two most widely researched directions.